Category Archives: WFU

Wide Field Upgrade is the commissioning effort to prepare for the new improved HET with the new secondary and instrument suite.

The week of March 7

We have had a great couple of weeks at the HET. Last week we brought out down the focal plane assembly to work on the field calibration unit and replace the acquisition camera. Both went well although some further work on a few specific calibration lamps and changes to the acquisition power supply will still be required. This work took the telescope down for a few days. The telescope control software also went through a large upgrade during this period. With improvements to guiding and offsetting. One of the more complicated aspects that has been commissioned is the ability of offset the telescope and the guide probes so that we can move from one instrument to another while keeping the same guide star.

The LRS2 commissioning team came back out to West Texas from Austin and despite very poor weather they were able to train a few of the resident astronomers in the use of the LRS2 instrument and the calibration scripts. When the skies did finally clear they were able to determine the positions of the LRS2-B and LRS2-R on the acquisition camera. The LRS2 team has departed but the instrument has been left online for the resident astronomers to gain familiarity and perhaps start characterizing on-sky.

The week of February 8

This week saw the return of the LRS2 but in its full glory, ie. both the B and R components. The instrument cooled down quickly and we were ahead of schedule by the third day. We did have a few problems with calibrations lamps, communications with the TCS, and inexperienced operators but all of that is expected when you get FIRST LIGHT with a science instrument on an effectively brand new telescope.

One of the highlights of the observing was moving back and forth between LRS2-B and LRS2-R, an observing mode we hoped to achieve eventually but were able to take advantage of in this first official commissioning run. The Austin team will be back in the beginning of March. In the mean time the LRS2 instruments remain at the telescope pumped down and cold to test the stability of the cryogenic system.

The week of January 25

Happy new year. We did conduct some engineering operations over the holidays. Most of this involved small changes to the pointing and collecting data to remove the remaining guider drift we continue to see. We made some changes to the lowest order of the pointing corrections.

We have also been able to move specific stars into the guide probes and wave front sensors. This past week was the first time we actually had stars in two guide probes and in the wavefront sensor at the same time.

At the beginning of January we began commissioning of the calibration wavefront sensor using the deployable wavefront sensor in the central IFU position. This is done with geo-stationary satellites and our procedure is to align the primary mirror and then setup on the deployable wavefront sensor, null out any tip, tilt and focus remaining and then image through the calibration wavefront sensor. Once we have done that a few times back and forth we offset the telescope so that we can image the satellite with the operation wavefront sensor 576″ away. We can now do this repeatably.

The day staff has been getting electronics setup in the 2nd VIRUS enclosure and dealing with small problems that have been popping up with the tracker and other sub-systems. The most critical of these was an uncontrolled warm up of the LRS-r. This was caused by a fault in the liquid nitrogen system and the normal warnings that normally go out were shunted into a dead phone line. The LRS-2 was sent back to Austin for diagnostics and new safeties will be put in place. We hope to have LRS-2 back out here next week.

We have hired a new Telescope Operator and he is being trained. We do have a open position for a computer system admin person currently posted so if you have anyone who wants to get away from the hustle and bustle of city life but still work with cutting edge equipment have them look at the UT job posting.

The week of Dec 7

Two weeks ago was the HET Board of Director’s meeting in Penn State. The meeting lasted two days and the Board got status reports from HET operations and each of the instrument teams. The main news is that that LRS2 is expecting to reach first science in the first few months of 2016, HRS2 is going to start commissioning before Summer 2016 and VIRUS units are going to be coming in over the next 9 months. The Board was encouraged by the progress being made and hopes that we can continue the pace. They were also pleased to see the progress being made on HPF and were impressed with the clean room facility tour they were given. No major changes or action items were reported by the Board.

Once the commissioning team returned from Happy Valley, we went right back to work and were able to push along the closure of one of our major metrology loops, the guide probes. I am pleased to report that we are able to guide at any telescope Az for full trajectories with the probes at any position within their range.

The week of Nov 16

This week we had lots of visitors from Austin. We had new equipment to install on the tracker, for example, the field calibration unit. This is the device at the entrance aperture of the corrector which can shine light from flat lamps or from line lamps for calibration of our spectrographs.

We also had delivery of the LRS2-R spectrograph and a temporary installation of a VIRUS unit. The VIRUS unit was put in place mostly to test the mechanics of the enclosure cooling and control systems. We were able to get our first spectrum with both spectrographs. By the end of the week we were able to find a guide star with guide probes at one specific orientation and then hold a science target on a specific location of the LRS2-R and VIRUS IFUs. We then could move the target around the IFU to confirm the direction within the IFU. Even the weather cooperated finally, with some near photometric conditions we put a spectrophotometric standard on the LRS2-R. On the VIRUS unit we confirmed that the dither mechanism does work and we were able to move a target from one fiber to the adjacent fiber.

We will be moving back to guide probe commissioning in the coming week and may not get back to LRS2 instrument commissioning until January.

All in all a very exhausting, but good week!!!

The week of Nov 9

This past week we had a large fraction of the Austin engineering and software groups visiting the telescope. This gave us an opportunity to make rapid improvements in the Telescope Control System and start getting the telescope ready for the installation of the first spectrograph.

With the removal of the Deployable Wave Front Sensors we were able to re-enable the guide probes and continue with their commissioning. Despite rather poor weather much of the week, the big milestone reached was the ability to now guide on the guide probes.

The software and engineering teams remain at the HET this next week and if we have some clear skies we will continue the push on commissioning of both the guide probes and refinement and characterization of the mount model.

The week of Nov 2

We have big news this week. We have officially declared first light on the wide field upgrade for the HET. Please see:

Upgrade HET Sees First Light

This is the first light image from the 29th of July.   The image quality is 1.3" and we were only an arcminute from the target of interest.

This is the first light image from the 29th of July. The image quality is 1.3″ and we were only an arcminute from the target of interest.

Now that verification of the WFU corrector optics is complete we can get started on operational aspects of the new HET. This next week we are collecting data with the guide probes which will allow us to start the software commissioning of the guide probes next week.

We have cleared out the temporary clean room we installed in our loading bay so we can now have access from the mirror lab to the HET dome floor. This means that the mirror team will start doing mirror re-coating in the near future.

Other interesting news comes from our big hail storm in the past month. It broke many car windows damaged a few of the residents’ roofs and below we show the damage to the HET dome. Thankfully no leaks that we can detect.

In the big hail storm of 2015 that fully covered the road in several inches of hail the HET dome took a bit of a beating.

In the big hail storm of 2015 that fully covered the road in several inches of hail the HET dome took a bit of a beating.

Since the last blog posting we had our most senior telescope operator retire. We wish her well in all of her next adventures. In the mean time we have posted the position and it is open now, probably for two more weeks.

Telescope Operator Position now Open

The week of Oct 5

Despite nearly a month of heavy smoke, fog, rain and hail we have completed our tests with the Deployable Wave Front Sensors (DWFS) on the Focal Plane Assembly (FPA). In these tests we were putting geostationary satellites on the various DWFS units across the FPA. While the final report on the image quality of the Corrector optics are still pending we do know that the image quality is good enough to move on to the next stage of commissioning. We will likely leave the central DWFS unit in place to allow us to check and re-optimize the image quality as we make further changes to the mount model. Below is an image of what a single bright star looks like through the DWFS unit. You can see that the light is broken up into small sub-aperatures (not each mirror) and then the distribution of those spots, ie. how evenly they are spaced, can tell you about the deformations of the optical system. The Corrector is the area in the center without the spots. The tracker rails run from 2 o’clock to 8 o’clock where the lower X work platform is still installed.

dwfs8_brightstar

On the staffing side of operations the good news is that we are fully staffed. We have a new electrical engineer and a optical technician starting this past week and a telescope operator joined us the week before. The bad news is that one of our other telescope operators has decided to retire as of the 14th of this month. We will be posting the position in the coming week.

The mirror team has begun an intensive inspection and repair regiment for the primary mirror actuators and electronics. They have found a number of damaged connectors and replacement of these connectors seems to have improved reliability considerably.

We have successfully tested the cooling of the first VIRUS enclosure. The cooling is accomplished by bringing in glycol that is only slightly cooler than the dome ambient temperature and then additional cooling from a small thermo-electric cooler is added when needed. This test was accomplished with no VIRUS units installed but artificial heat loads induced. We will have to wait to see how the system performs under the full load of the VIRUS units.

The week of August 24

While much work has continued during the day focus has begun to shift to support of night operations and commissioning. The blogger has moved back to working the night shift so updates on day operations may be second hand.

The day staff have installed the acquisition camera, pupil viewer, guide probes and Deployable Wave Front Sensors (DWFS) on the Focal Plane Assembly (FPA). In addition, they have gotten glycol lines hooked up to the VIRUS enclosures and those lines have been pressure tested. The glycol lines to the top of the telescope for cooling the electronics boxes and cameras are still on order.

The software team in Austin has been out to the telescope twice over the past few weeks to commission various aspects of the telescope and mount model. The preliminary mount model has been completed and we can now track on a star and keep constant focus and tracker induced coma through a trajectory. The DWFS will be used in the next weeks to determine the exact tip and tilt zero point to minimize the tracker induced coma and to search for any problems with the mounting of the tracker optics. This will be done on geostationary satellites which allow us to have a constant point source at infinity for long periods of time with minimal tracker motion. We have confirmed that we can acquire these geostationary satellites on the Acquisition camera.

The night staff is also in the process of commissioning the guide probes. We have been able to obtain star fields in both guide probes and the acquisition camera at the same time. With this data and probably a few more similar data sets at different guide probe positions we will be able to set the zero points to pointing with the guide probes.

While we continue to work out small problems and inefficiencies with the tracker electronics and telescope control software overall the telescope is coming along well and some ways is far more impressive than the old HET.

The week of July 20

This week the visiting engineers from Austin aligned the Video Alignment Telescope (VAT) to the corrector optics and then aligned the corrector to the the center of the primary mirror. This sets our axis for making all trajectories. We will now set the distance from the primary to the VAT. Work will continue over the weekend and we will then turn the VAT around and measure the angle from the VAT up to the CCAS instrument. This last will be night work. So far things are progressing on schedule.